### v1

# Effects of lower floating-point precision on scale-resolving numerical simulations of turbulence

Abstract
Modern computing clusters offer specialized hardware for reduced-precision arithmetic that can speed up the time to solution significantly. This is possible due to a decrease in data movement, as well as the ability to perform arithmetic operations at a faster rate. However, for high-fidelity simulations of turbulence, such as direct and large-eddy simulation, the impact of reduced precision on the computed solution and the resulting uncertainty across flow solvers and different flow cases have not been explored in detail and limits the optimal utilization of new high-performance computing systems. In this work, the effect of reduced precision is studied using four diverse computational fluid dynamics (CFD) solvers (two incompressible, Neko and Simson, and two compressible, PadeLibs and SSDC) using four test cases: turbulent channel flow at $Re_{\tau}=550$ and higher, forced transition in a channel, flow over a cylinder at $Re_{D}=3900$, and compressible flow over a wing section at $Re_{c}=50000$. We observe that the flow physics are remarkably robust with respect to reduction in lower floating-point precision, and that often other forms of uncertainty, due to for example time averaging, often have a much larger impact on the computed result. Our results indicate that different terms in the Navier–Stokes equations can be computed to a lower floating-point accuracy without affecting the results. In particular, standard IEEE single precision can be used effectively for the entirety of the simulation, showing no significant discrepancies from double-precision results across the solvers and cases considered. Potential pitfalls are also discussed.
keywords:
Computational Fluid Dynamics, turbulence, direct numerical simulation, floating-point precision
1 Introduction
Computational fluid dynamics (CFD) has become an essential tool in both academic research and industry, spanning a diverse array of applications. Over time, a variety of models and numerical methods have been developed to integrate the governing equations of fluid motion, each tailored to specific use cases, desired accuracy, and computational constraints. Yet, until rather recently, all of them typically relied on IEEE double precision floating point numbers (FP64) to numerically compute the solution.
Due to recent shifts in hardware manufacturing tailored to decrease the energy consumption of floating-point operations and drive up performance for artificial intelligence (AI) applications, hardware support for lower precision floating-point numbers has become increasingly prevalent [1]. On these new platforms, lower precision offers both higher performance, higher energy efficiency, as well as a smaller memory footprint, reducing the amount of data movement necessary. All of these developments in the end lead to significant savings in both time and energy and thus provide large monetary savings for large-scale computations. As CFD practitioners, where most codes are limited by memory bandwidth due to large relatively sparse linear algebra systems, the use of FP32 instead of FP64 would, for example, move the bandwidth roofline by a factor 2, indicating an up to 2X of performance would be attainable for a bandwidth limited code [2]. A simple roofline comparison between two relatively common GPUs, Nvidia A100 and Nvidia GeForce RTX4080, is shown in Figure 1. The simple performance model clearly illustrates how the roofline for the RTX4080, with a significantly lower performance for FP64, improves for FP32 and is comparable to the more expensive A100. Considering that the operational intensity $I$ for a given code also improves with lower precision, this points to the significant gains that can be enabled by lower precision. In practice, a factor of 2×2$\times$2 × in performance is rarely achieved due to other factors such as kernel launch latencies or communication overhead in parallel communications.
Figure 1: Roofline for the Nvidia A100 and Nvidia GeForce RTX4080 in double (FP64) and single (FP32) precision. The solid lines represent the roofline (the maximum attainable performance) for the two architectures as a function of operational intensity $I=\pi/\beta$, defined as the fraction between the peak performance $\pi$ and the memory bandwidth $\beta$ of the computing unit. The dashed lines represent the peak performance $\pi$ for the two architectures and the dotted line the performance limit based on the time needed to load data from memory, $\beta I$. Most CFD codes today operate in the domain limited by $\beta I$.
While there are clear opportunities for performance gains, an important question is to what extent these lower floating point formats can be used for scale-resolving simulation of turbulence without sacrificing the required accuracy. This is what we aim to assess in this work.
This question is under active investigation by the research community and can be viewed as part of a general trend across many disciplines [3]. A full review of works on lower-precision arithmetics and CFD is outside the scope of this article, but a few selected references are provided for the benefit of the reader. The first example is from simulations of weather and climate. Several articles from Klöwer, Paxton, Düben, and Palmer et al. [4, 5, 6] highlight that lower precision and less exact computing open up for large performance improvements. In the realm of compressible flows, the CFD solver PyFR, [7, 8], used single precision (FP32) in the context of a high-order flux-reconstruction scheme for implicit large eddy simulation for a wide range of cases. Methods that enable mixed-precision computing while maintaining acceptable levels of accuracy have been proposed and investigated, for instance, in the context of the finite volume solver OpenFOAM [9], further encouraging results using OpenFOAM were also found by Brogi et al. [10]. Haridas et al. [11] explored the possibility of using neural networks to correct errors introduced using reduced precision arithmetic in simulating fluid dynamic problems. The feasibility of conducting mixed-precision operations in the context of high-order compact schemes were investigated by Song et al. [12] using the PadeLibs code for CFD. Wang et al. [13] proposed a mixed precision strategy in the finite volume method for unstructured grids that used high precision near solid bodies and lower precision far away from them. Freytag et al. [14] studied the performance and power efficiency of using reduced- and mixed-precision arithmetic for CFD. Bhola and Duraisamy [15, 16] performed analysis of errors incurred due to rounding in mixed precision computations. Walden et al. [17] investigated the speedup obtained on GPUs with the FUN3D CFD code by using reduced-precision arithmetic for their memory-bound linear solver kernel. Grout [18] used reduced-precision for the low-order time integration methods that were used to build higher-order methods in the so-called spectral differed correction method used in combustion CFD and studied the resulting rate of convergence. In order to reduce the communication bottleneck, Rogowski et al. [19] investigated reduced precision for communication in the context of Discontinuous Galerkin method using the SSDC CFD code. Finally, for CFD based on the Lattice Boltzmann method rather than solving the Navier-Stokes (NS) equations, [20] demonstrates the possibility to accelerate the solver using lower precision. It is also worth noting that finite volume codes used in industry, both open-source and commercial, offered the possibility to compile in FP32 long before the recent increase in low-precision computer hardware. There are also many mixed-precision algorithms for important linear algebra operations used in CFD, such as preconditioners and iterative solvers, to obtain a result in full double precision [21, 22, 3, 23]. This potentially suggests a wider scale of adoption of FP32 than what is documented in the scientific literature.
In spite of these success stories, it is important to note that due to the nonlinear nature of the NS equations, cases where numerical precision will have a larger-than-expected impact can be found. In particular, in the context of dynamical systems, if there exist several attractors of the flow or a symmetry that is sensitive to small disturbances, lower floating-point precision can be detrimental to the validity of the simulation [24, 25]. However, if the simulation only has one attractor, and the results are insensitive to small thermal fluctuations, which in many ways can act similarly to low-precision [26], lower precision might be utilized effectively.
The question in the focus of this article is whether reduced precision arithmetic can be successfully employed in direct numerical simulations (DNS) and wall-resolved large-eddy simulation (LES). The impression of the authors is that the DNS community is generally skeptical towards abandoning FP64. There are good reasons behind that. DNS are usually conducted to obtain the ground truth for a given flow and study all possible subtleties in its behavior. Therefore, even small errors are considered unacceptable. Furthermore, a new DNS is usually conducted at the limit of the computing budget, and simply trying out a low-precision run can be perceived as not worth the risk.
Nevertheless, the main outcome of the DNS is often primarily expressed in a statistical description of the flow, and an argument can be made that it must, to a degree, be robust to small numerical errors, the origin of which may be both discretization and the precision of arithmetics. Indeed, there are already documented successful attempts in using FP32 for DNS of homogeneous isotropic turbulence (HIT) [27, 28, 29, 30], with no discernible differences observed between the FP32 and FP64 runs. While this is very encouraging, to trigger a shift across the wider high-fidelity CFD community, it is necessary for similar evidence to emerge across a broader range of flow cases and numerical techniques. The goal of the current work is to be a step in that direction, and toward a deeper understanding of the potential impact of lower floating precision on high-fidelity CFD. This builds on some preliminary work performed by Karp et al. [31, 32, 33].
To that end, our study makes use of four different solvers (Neko, SSDC, Simson, PadeLibs), that differ in both the formulation of the governing equations and the approach to discretization. We consider four use cases, with a focus on wall-bounded flows mostly in the turbulent regime. The latter is motivated partially by the scientific interests of the authors, but also by the fact that turbulence near the wall has a particularly rich and complex structure that could potentially be disrupted due to floating point errors. The four use cases cover several important flow classes: internal flows (the channel flow test case), transition to turbulence (the Tollmien-Schlichting wave test case), external flows over bluff bodies (flow around a cylinder), and external flows with influence of compressibility (the NACA-0012 case).
In our analysis, we go below FP32 and also consider precision that is not necessarily implemented in current hardware. To facilitate that, we rely on a software emulation of the precision. However, we have also developed native FP32 versions of Neko, Simson, and, partially, SSDC. This work reveals that naively reducing the precision of all reals may lead to various pitfalls. Therefore, besides presenting simulation results, we also summarize practical experiences in adapting our codes to FP32, in order to aid other members of the community in similar efforts.
The paper is structured as follows. Section 2 provides a description of the common floating-point formats. Section 3 introduces the four CFD codes that will be used for the present study. The remainder of the paper presents the impact of precision on four different flow cases: Turbulent channel flow will be analysed in Section 4, and Section 5 treats transitional channel flow. External flow with separation is discussed in Section 6, followed by the flow around wings in Section 7. Practical experiences and conclusions wrap up the paper in Sections 8 and 9.
2 Floating-point numbers
In this work, we are interested in the impact of floating point formats and their ability to represent relevant physics in a turbulent flow simulation. A floating-point number is defined by a number of mantissa bits $b$ and exponent bits $b_{e}$ together with one sign bit $s$ dictating the sign of the floating-point number. If we let $e$ be the value of the exponent (as an unsigned $b_{e}$-bit integer), and $c_{i}$ be the $i$th least significant bit of the mantissa, the value for a given normal floating-point number is
$(-1)^{s}\left(1+\sum_{i=1}^{b}c_{b-i}2^{-i}\right)\times 2^{e-(2^{b_{e}-1}-1)}.$ (1)
We consider floating-point formats between 8 and 64 bits as listed in Table 1. In addition to the normal floating-point numbers, there are also special numbers such as $\pm$ infinity, not a number (NaN) and subnormal numbers.
Name bits $b$ $b_{e}$ $\varepsilon=2^{-b-1}$
FP64 64 52 11 $2^{-53}\approx 2\cdot 10^{-16}$
FP32 32 23 8 $2^{-24}\approx 6\cdot 10^{-8}$
FP16 16 10 5 $2^{-11}\approx 5\cdot 10^{-4}$
bfloat16 16 7 8 $2^{-8}\approx 4\cdot 10^{-3}$
E4M3 8 3 4 $0.06252^{-4}=$ 0.0625
E5M2 8 2 5 $0.1252^{-3}=$ 0.125
Table 1: Different floating-point formats. The rounding machine epsilon $\varepsilon$, $|u_{\text{FP}}-u|<\varepsilon u$ for some real number $u$, is the largest round-off error introduced due to floating-point precision.
Floating-point numbers have a great strength in their large range and use of a relative rounding error, compared to an absolute one such as for fixed-point formats. This means that when operations are carried out on numbers of comparable amplitude the effect of rounding error is relatively small. However, issues can, for example, arise when performing computations with two numbers $x$ and $y$ where $x\ll y$. In this case, if $x<\varepsilon y$, the computation may be subject to so-called stagnation. A typical example is when performing summations of long arrays and an individual element in the array is smaller than the total sum, which can yield a significant rounding error. This would, for example, yield inaccurate dot products, large errors after many time steps as well as possibly impact the collection of statistical quantities.
This aspect of stagnation is a general issue in dynamical systems as well, and efforts to avoid it through approaches such as stochastic rounding have been suggested [6]. However, floating-point numbers with deterministic rounding (rounding to the nearest) are the most readily available and commonly used in modern computing systems, and in this work we limit ourselves to this type of quantization.
The issue of when the machine epsilon plays a large role for CFD, can be clearly seen through inspection of the NS equations as well. If we consider the non-dimensionalized incompressible NS equations,
$\begin{split}\nabla\cdot\mathbf{v}&=0,\\ \frac{\partial\mathbf{v}}{\partial t}+(\mathbf{v}\cdot\nabla)\mathbf{v}&=-\nabla p+\frac{1}{Re}\nabla^{2}\mathbf{v}+\mathbf{F},\end{split}$ (2)
where $\mathbf{v}$ is the velocity field, $p$ the pressure, $\mathbf{F}$ an external forcing and $Re=U/\nu L$ is the Reynolds number defined for some suitable characteristic velocity $U$ and length scale $L$ and the kinematic viscosity of the fluid $\nu$. As can be seen, the choice of non-dimensionalization and the $Re$ number immediately provides a connection between the strength of the different terms of the equations. Taking the extreme case when for example $\frac{1}{Re}<\varepsilon$, the time integration of the system would be significantly affected by numerical round-off.
If we consider $\mathbf{u}^{i}$ to be the state of the flow at time step $i$ and time $i\Delta t$. It is a discretized vector of length $n$ $\mathbf{u}\in FP^{n}$ where $FP$ is the set of floating point numbers possible to be represented for a floating-point format FP. The total simulation up to time $i\Delta t$ can then be described as the set of realizations of the flow $\mathbf{U}=\{u^{0},\ldots,u^{i}\}$, where $\mathbf{U}$ is constructed through some map from state $\mathbf{u}^{i}$ to $\mathbf{u}^{i+1}$
$\mathbf{u}^{i+1}=f(\mathbf{u}^{i}).$ (3)
In this investigation we make an attempt to see how $f$, which is described by the numerical method, is affected, regardless of and the elements in $\mathbf{u}$ are not the focus, but rather how different terms in the Navier-Stokes equations and the state vector $\mathbf{u}$ are affected by lower floating point precision. This view of a numerical simulation holds for any numerical discretization of the flow. Evaluating different numerical methods such as finite volume, finite elements or similar would in this model equal the choice of the element in the state vector and how the map $f$ is computed.
3 Methodology
Throughout this work, we consider floating-point precision as a rounding operation from some state $u$ to a rounded state $\tilde{u}$. In our work, for all solvers considered in the following sections, all floating-point numbers below FP32 are emulated with CPFloat [34]. Several opportunities to incorporate rounding into the governing equations exist, and the question is how different terms, such as the advective non-linear term depend on the numerical precision. The nonlinear term is especially important for turbulent cases where $Re$ is large, the equations are chaotic and small disturbances make two trajectories diverge quickly. As such, analytical or deterministic approaches to assess the accuracy of the simulation are no longer applicable. In addition, in a numerical setting where the equations are discretized, the interaction and dependence on an accurate numerical format can be drastic. As such, we consider several different cases to evaluate if a simulation can be run entirely in lower precision, as well as only perturbing specific terms in the governing equations. In particular, we distinguish three approaches:
  1. 1.
    Full FP32. The entire solver is run using IEEE single precision. This is the only case where the lower precision is not emulated in this work.
  2. 2.
    State rounding. Casting $\boldsymbol{u}^{i}$ in lower precision, while the solver operates in FP64. Simulating the perturbed system, $\mathbf{u}^{i+1}=f(\tilde{\mathbf{u}^{i}})$, where the state is constrained to a lower floating point precision at each time $i$.
  3. 3.
    Term rounding. Different terms in the NS equations, such as the convective or viscous term or both, are represented in lower precision; for example, for the convective term in the incompressible formulation would be computed according to $\widetilde{(\mathbf{v}\cdot\nabla)\mathbf{v}}$ with the rounding operator.
In order to assess how these perturbations impact different numerical schemes and different formulations of the Navier–Stokes equations, we consider four different flow solvers with different discretizations and characteristics. In light of these differences, the specific way the rounding is applied differs somewhat between the codes. These details are discussed for each solver individually.
3.1 Software and numerical methods
3.1.1 Neko
Neko is based on a continuous Galerkin spectral-element framework with a special focus on the incompressible Navier–Stokes equations, with extensive support for heterogeneous computer architectures [35]. The code has excellent scaling demonstrated to thousands of GPUs and was nominated for the Gordon Bell Prize in 2023 [36]. The solver uses high-order hexahedral spectral elements (polynomial order 7 for the tests here), with the $P_{N}-P_{N}$ method for velocity–pressure decoupling, a third-order semi-implicit time integration method, and dealiasing of the convective term using the $3/2$-rule [37, 38]. The following tests are performed with Neko: Full FP32 (representing all floating-point numbers and executing operations in FP32), perturbation of the convective term (denoted Convective/Conv. FPX in tables and plots for precision FPX), and state rounding (denoted State FPX). Note that for the Full FP32 runs, the mesh files used in the simulations were still the same, but was directly converted to FP32 when loaded into Neko.
3.1.2 Simson
Simson [39] is a fully spectral code for channel and boundary-layer configurations, based on Fourier discretization in the streamwise and spanwise directions, and Chebyshev expansion in the vertical (wall-normal) direction. The mesh is equidistant in the wall-parallel directions, and follows a Gauss–Lobatto distribution in the wall-normal direction. Standard dealiasing using the $3/2$ rule is performed in the Fourier directions only. All solvers are direct in velocity–vorticity formulation; thus, no tolerances need to be specified. The classical fourth-order Runge-Kutta (RK4) method is used for time integration. Tests performed using Simson include Full FP32, Convective FPX, and State FPX. Convective FPX is implemented in Simson by rounding of the convective term after its calculation. State rounding is implemented by rounding of all relevant fields (velocities and vorticities) at the start of each RK4 substep. Since precisions lower than FP32 are emulated, in both State and Convective FPX the rest of the operations are performed in FP64.
3.1.3 SSDC
SSDC implements a high-order entropy-stable discontinuous collocated Galerkin method for the compressible Navier–Stokes equations [40]. The SSDC framework is built on top of the highly scalable Portable and Extensible Toolkit for Scientific Computing (PETSc) [41], its mesh topology abstraction (DMPlex), and its scalable differential–algebraic equation solver components. The spatial discretization features $hp$-adaptive capabilities on unstructured quadrilateral/hexahedral meshes. Support for nonconforming meshes relies on the p4est software library [42] and its bridge to PETSc’s DMPlex. Triangle/tetrahedral meshes are converted on the fly into quadrilateral/hexahedral elements; uniform and non-uniform mesh refinement algorithms are also available. The collocation nodes inside each element are distributed according to the Gauss–Legendre–Lobatto quadrature points. The solver has demonstrated good strong parallel scaling up to at least 700 000 CPU cores on Shaheen III supercomputer hosted at KAUST. The time integration is explicit and performed using the Runge–Kutta scheme of Bogacki–Shampine [43] of order three with four stages with the first-same-as-last property. This Runge–Kutta scheme has an embedded second-order method used to implement adaptive step size.
Full FP32 capabilities were partially implemented in SSDC as part of this work, currently limited to precomputed metric terms in FP64 on top of a fully FP32 computation (see Section 8 for details). State rounding was not considered. Regarding term-rounding, the explicit time integration implies that the rounding is applied to the corresponding term in the right-hand side, at each stage of the Runge–Kutta scheme. We implemented the possibility to treat the viscous and convective operators separately. Moreover, the rounding can be applied either on the fields before evaluating the operator, on the evaluated operator’s result, or on both. To be aligned with the notation for the other codes, Convective FPX refers to rounding the convective operator’s output. Applying the rounding to both terms, prior and posterior to the application of the respective operators is referred to as Combined FPX.
3.1.4 PadeLibs
PadeLibs is a Navier–Stokes solver for high-resolution simulations of compressible turbulent flows [44]. The numerical discretization uses sixth-order compact finite-difference methods with collocated variable storage and staggered flux assembly. The simulation framework used in PadeLibs is robust to aliasing errors and has high accuracy in resolving diffusive fluxes at small scales. In this work, round-off effects are investigated by rounding the convective (inviscid) fluxes to a precision FPX (Convective FPX) after they are assembled before taking the divergence operations. The rounded results still keep the double-precision format (FP64), although the emulated round-off errors are introduced. All the differential and interpolation operations are consistently calculated in double-precision format. The operator coefficients are all at double-precision accuracy, and the round-off errors are added only from the input. For the incompressible test cases, the Mach number is set to be $0.25$.
4 Fully developed turbulence
The first test case is turbulent channel flow at $Re_{\tau}=550$ in a relatively modest domain of $2\pi\delta\times 2\delta\times\pi\delta$. Here, $\delta$ is the half-height of the channel and $Re_{\tau}$ is the friction Reynolds number. The resolutions follow standard practice for high-order simulations of wall turbulence: $\Delta x^{+}\approx 12$, $\Delta z^{+}\approx 5$, and $\Delta y^{+}$ similar to, e.g. , [45]. In particular, for SSDC and Neko, the first off-wall node is located at $y^{+}\approx 0.45$. Table 2 summarizes the different successful simulations and their corresponding simulation parameters. The maximum difference of the first- and second-order moments for the streamwise velocity, $u$, in inner units are compared with that of FP64 run using the respective code as well as with the DNS of Lee and Moser (LM) [46]. These error measures are denoted using the expression shown in Equation (4), where $q$ can be $u$ or $u^{\prime}u^{\prime}$, and ``ref"``$ref"$ is replaced either with “FP64” if the FP64 run (from the corresponding CFD code) is used, or is replaced with “LM” if the data from LM [46] is used instead.
$\mathcal{E}_{q_{\rm ref}}=\text{max}\left(|\langle q\rangle^{+}-\langle q\rangle^{+}_{\rm ref}|\right)$ (4)
Setup $Re_{\tau}$ Avg. time $\mathcal{E}_{U_{\rm FP64}}$ $\mathcal{E}_{u^{\prime}u^{\prime}_{\rm FP64}}$ $\mathcal{E}_{U_{\rm LM}}$ $\mathcal{E}_{u^{\prime}u^{\prime}_{\rm LM}}$
Neko
Full FP64 548 $66.6\delta/u_{\tau}$ 0.008 0.125
Full FP32 557 $150.0\delta/u_{\tau}$ 0.179 0.126 0.154 0.107
Convective FP32 550 $105.5\delta/u_{\tau}$ 0.109 0.103 0.090 0.112
State FP32 549 $93.4\delta/u_{\tau}$ 0.204 0.172 0.185 0.115
Convective FP16 550 $84.8\delta/u_{\tau}$ 0.118 0.103 0.098 0.111
State FP16 683 $42.5\delta/u_{\tau}$ 3.797 2.594 3.847 2.073
Convective E5M2 552 $49.9\delta/u_{\tau}$ 0.110 0.128 0.084 0.138
Convective E4M3 549 $53.1\delta/u_{\tau}$ 0.229 0.132 0.210 0.136
Simson
Full FP64 546 $43.6\delta/u_{\tau}$ 0.16 0.14
Full FP32 546 $43.6\delta/u_{\tau}$ 0.09 0.10 0.07 0.10
Convective FP16 544 $43.5\delta/u_{\tau}$ 0.16 0.18 0.02 0.12
State FP16 543 $43.4\delta/u_{\tau}$ 0.23 0.13 0.07 0.05
Convective E5M2 541 $43.3\delta/u_{\tau}$ 0.35 0.24 0.20 0.27
State E5M2 337 $26.9\delta/u_{\tau}$ 20.8 7.5 20.8 7.6
Convective E4M3 300 $14.0\delta/u_{\tau}$ 20.2 36.2 20.2 36.2
State E4M3 1284 $7.9\delta/u_{\tau}$ 15.6 34.5 15.6 33.9
SSDC
Full FP64 549 $43.95\delta/u_{\tau}$ 0.033 0.077
Combined FP32 550 $43.96\delta/u_{\tau}$ 0.054 0.082 0.042 0.113
Combined FP16 549 $43.93\delta/u_{\tau}$ 0.059 0.048 0.086 0.077
Convective FP32 549 $43.93\delta/u_{\tau}$ 0.028 0.104 0.024 0.114
Convective FP16 549 $43.90\delta/u_{\tau}$ 0.024 0.051 0.045 0.081
Table 2: Details for the different channel-flow simulations conducted in a domain of size $2\pi\delta\times 2\delta\times\pi\delta$. The reported error values (computed using Equation 4) are in inner units and might be impacted by the averaging times. The compressible codes (SSDC and PadeLibs) are expected to have higher error levels compared to reference data [46] due to compressibility effects. Additional tests were carried out using Simson, including the effect of Reynolds number, resolution, domain size, and time step which are shown in Table 3.
As shown in Table 2, the observations from the two sets of error measures are quite similar, hinting at the confidence in the simulations that were performed. As a side note, the slightly higher errors observed for Full FP32 in Neko compared to the cases following right after it in Table 2 could be due to the extra errors introduced when writing out the statistics in single precision using the csv format (Full FP32 was the only case where the file was written out using single precision).
Figure 2: Mean streamwise velocity profiles of the turbulent channel flow simulated using different precisions at $Re_{\tau}=550$ using Neko, SSDC and Simson. All curves agree reasonably well with one another and to the data from Lee & Moser [46] (not shown here), except State FP16 from Neko; and State E5M2, Conv. E4M3, State E4M3 from Simson. Different codes are shown using different line styles (as shown in the legend on the left) and the different roundings are shown using different colors (as shown in the legend on the right). All these cases are also compared against each other in Table 2.
Figure 3: Root-mean-square of velocity fluctuations from the turbulent channel flow simulated using different precisions at $Re_{\tau}=550$ using Neko and SSDC. All curves agree reasonably well with one another and to the DNS data from Lee & Moser [46] (not shown here), except State FP16 from Neko; and State E5M2, Conv. E4M3, State E4M3 from Simson. Different codes are shown using different line styles (as shown in the legend on the left) and the different roundings are shown using different colors (as shown in the legend on the right). All these cases are also compared against each other in Table 2.
Overall, the results of the different simulations were largely unaffected by low precision down to FP16. Especially for first-order moments the solution is not visibly sensitive. This is shown in Figure 2 where the results from Neko, Simson and SSDC are compared against each other. Note that the Neko, Simson and SSDC cases that were named “Full FPXX” or “Combined FPXX” (as seen in Table 2) are named in Figures 2 and 3 as just “FPXX” for the sake of brevity. This was done since they are very close in their rounded representation and indistinguishably similar in these plots. The mean streamwise velocity profiles, shown in Figure 2, that are notably different are that of the state rounding to FP16 in Neko; and convective rounding to E4M3, state rounding to E4M3 and E5M2 in Simson. For the second-order moments shown in Figure 3, all cases show excellent agreement with FP64 as well as with the DNS data from Lee & Moser [46] (not shown in figure, but shown in Table 2), except the same cases mentioned above for the mean.
The lowest precision that worked fine (i.e., giving correct profiles) for the different codes are as follows: for Neko, Convective FP16 (but not State FP16); for Simson, Convective and State FP16; for SSDC, Combined and Convective FP16. Combining these results FP16 seems to be the edge case, however no code was run entirely in FP16 (the closest being SSDC that was run using Combined FP16). It should be noted that Table 2 and Figures  2 and  3 show cases that did not diverge. The entire set of runs that were carried out can be found in the Table 7. This means that SIMSON was the code that was most resilient to reduced precision, for not having diverged down to E5M2 (although it resulted in wrong profiles), which is probably due to the Fourier formulation of the code.
4.1 Additional assessments
The importance of arithmetic precision was further investigated for parameters such as the Reynolds number, resolution, and domain size, and for higher-order moments and more complex statistics such as the budget terms.
The impact of the number of modes in the streamwise and spanwise directions were tested using Simson by increasing the resolution for a fixed domain size, up to a resolution of $\Delta x^{+}\approx 9.0$ and $\Delta z^{+}\approx 4.5$, as well as increasing the domain size to $8\pi\delta\times 2\delta\times 3\pi\delta$ for the higher resolutions (see Table 3). The impact of wall resolution was also tested by decreasing the number of Chebyshev modes in the wall-normal direction (not included in Table 3). These tests were only done in FP64 and FP32 and showed no statistically significant variation in statistics such as the mean velocity or Reynolds stresses.
Precision $Re_{\tau}$ Avg. time Domain size Resolution $\mathcal{E}_{U_{\rm LM}}$ $\mathcal{E}_{u^{\prime}u^{\prime}_{\rm LM}}$
Full FP64 544 $145\delta/u_{\tau}$ $8\pi\delta\times 2\delta\times 3\pi\delta$ $(9.0,0.04,4.5)$ 0.018 0.015
Full FP32 543 $145\delta/u_{\tau}$ $8\pi\delta\times 2\delta\times 3\pi\delta$ $(9.0,0.04,4.5)$ 0.019 0.013
Full FP32 999 $92\delta/u_{\tau}$ $8\pi\delta\times 2\delta\times 3\pi\delta$ $(9.8,0.03,4.6)$ 0.10 0.35
Full FP32 998 $18.8\delta/u_{\tau}$ $4\pi\delta\times 2\delta\times 1.5\pi\delta$ $(9.8,0.03,4.6)$ 0.12 0.11
Full FP32 998 $51.7\delta/u_{\tau}$ $2\pi\delta\times 2\delta\times\pi\delta$ $(9.8,0.03,4.6)$ 0.13 0.11
Full FP32 1985 $33.4\delta/u_{\tau}$ $2\pi\delta\times 2\delta\times\pi\delta$ $(10.9,0.04,6.5)$ 0.16 0.09
Table 3: Additional simulations of channel flow carried out using Simson. Resolution are reported in friction units as ($\Delta x^{+},\Delta y^{+}_{1},\Delta z^{+}$) for the streamwise, wall-normal (next to the wall), and spanwise resolutions.
The impact of Reynolds number (scale separation) was tested by conducting a simulation using Simson compiled with FP32 at the higher Reynolds number of $Re_{\tau}\approx 1000$ in a domain of size $8\pi\delta\times 2\delta\times 3\pi\delta$ (Table 3). The overall behavior was extremely similar to $Re_{\tau}\approx 550$ with no outstanding difference between FP32 and FP64 for the mean velocity and Reynolds stresses.
A more detailed analysis of budget terms in the transport equation of Reynolds stresses was also performed. Interestingly, the components related to pressure-velocity coupling (pressure-strain and pressure transport terms) were the only ones sensitive to the arithmetic precision at FP32, as shown in Figure 4. The behavior was similar in both $Re_{\tau}\approx 550$ and $Re_{\tau}\approx 1000$. However, we should note that since Simson uses the velocity–vorticity formulation, the instantaneous pressure does not enter the evolution of the flow and is computed as a separate step only if needed. Since the budget terms related to velocity gradients were robust to precision, we hypothesized that the observed differences were likely caused by a sensitivity to precision in the Poisson solver. This hypothesis was confirmed by recompiling the code in FP64, restarting the $Re_{\tau}\approx 1000$ simulation from snapshots written in FP32, and taking one time step only to recompute the pressure in FP64. This is shown in Fig. 4. Note that there is still a small difference (around 5%) in both pressure-related terms in the region $y^{+}\leq 10$ which is not observed in FP64. Therefore, additional assessments might be necessary before adopting FP32 for producing reference DNS datasets; for example, to assess whether the correct values can be recovered by allowing a few time steps in FP64 (equivalent to switching to FP64 during runtime and before outputting the fields).
(a) (b)
Figure 4: Budget terms in the transport equation of $\langle u^{\prime}u^{\prime}\rangle$ (a) and $\langle u^{\prime}v^{\prime}\rangle$ (b) for the turbulent channel flow at $Re_{\tau}\approx 1000$ using full FP32 (beige) and recalculated using one time step in FP64 (dark red) compared to the reference data from Lee & Moser [46] (dotted blue). Simulations are performed using Simson. Triangles and squares denote the pressure-strain and pressure transport terms, respectively.
To assess the sensitivity of the conclusions to the specific implementation in Simson, this test was repeated using Neko for the turbulent pipe flow at $Re_{\tau}\approx 1000$ (not shown here) in a domain of length $L_{z}=4\pi R$ (where $R$ is the pipe radius) and resolutions of $(\Delta z^{+},\Delta R^{+},(R\Delta\theta)^{+})\approx(5.3,0.5\rightarrow 10,5.3\rightarrow 10)$ in the streamwise, wall-normal, and azimuthal directions, respectively. The choice of pipe flow was motivated by having more complex mappings between the physical and computational space. Interestingly, all budget terms (not shown) were extremely similar for FP32 and FP64 and matched the reference data of Yao et al. [47] with no issues observed with the pressure-strain or the pressure transport term. This, in fact, reinforces the previous hypothesis that the observed issues are specific to the formulation and implementation in Simson, and may not be observed, or at least can be largely avoided, in other solvers.
In addition to the budget terms, higher moments of the solution, such as the third and fourth moments, can be impacted by the precision. This is illustrated in Fig. 5 for velocity skewness (i.e., $\langle u^{\prime 3}_{i}\rangle/\langle u^{\prime 2}_{i}\rangle^{3/2}$) and kurtosis (i.e., $\langle u^{\prime 4}_{i}\rangle/\langle u^{\prime 2}_{i}\rangle^{2}$), where clear fluctuations can be observed for $\langle u^{\prime 4}_{1}\rangle/\langle u^{\prime 2}_{1}\rangle^{2}$ in the region $y^{+}\geq 100$. This was found to be caused by two somewhat independent issues: (i) precision used for post-processing of data, and (ii) precision used for calculation and writing of the velocity fields. The impact of (i) was tested by performing the entire post-processing in FP32 for a simulation done entirely in FP64, where it was observed that similar oscillations still occurred. Interestingly, while item (ii) seemed more serious at first, its impact could still be removed by a method similar to what was done for the budget terms, i.e., by restarting the simulation using FP64, taking one time step, and rewriting the fields. This procedure completely removed the oscillations, as can be observed in Fig. 5, with values that were, within statistical significance, identical to a simulation carried out and post-processed in FP64. These observations, combined with the absence of such oscillations from the wall-normal and spanwise components of velocity (which do not have $\mathcal{O}(U_{b})$ mean components, leading to $\mathcal{O}(\varepsilon U_{b})$ errors), suggest that the sensitivity is most likely the result of a combination of the precision at which the velocity was written and the solver that calculates the streamwise velocity from wall-normal velocity and vorticity (as is done in velocity-vorticity formulation).
(a) (b)
Figure 5: Skewness (a) and flatness (b) of velocity components in the streamwise ($u_{1}$), wall-normal ($u_{2}$), and spanwise ($u_{3}$) directions in the turbulent channel flow at $Re_{\tau}\approx 550$. Colors from light to dark show cases that were run and post-processed in FP32, run in FP32 but restarted and post-processed in FP64, and run and post-processed in FP64. Simulations were performed using Simson with velocity-vorticity formulation. Similar behavior was observed at $Re_{\tau}\approx 1000$, with larger fluctuations.
Both the budget terms and higher velocity moments were tested for the turbulent channel flow at $Re_{\tau}\approx 2000$ (Table 3). The observations were nearly identical to $Re_{\tau}\approx 550$ and $Re_{\tau}\approx 1000$, except for the increased sensitivity of the moments with increased Reynolds number. We did not test whether the correct pressure-related budget terms or the third and fourth moments could be reconstructed by one small iteration in FP64. We also performed a turbulent pipe flow simulation at $Re_{\tau}\approx 2000$ in Full FP32 using Neko. Similarly to $Re_{\tau}\approx 1000$, the budget terms (not shown here) were indistinguishable from the reference data of Yao et al. [47], again hinting at the higher sensitivity of Simson to arithmetic precision.
One should note that the results presented here do not guarantee a similar behavior for significantly higher Reynolds numbers (such as $Re_{\tau}\approx 10,000$ or higher), especially when generating reference quality data for the community. However, we feel confident that for the majority of the simulations performed nowadays, e.g., for Reynolds numbers up to $Re_{\tau}\approx 2000$, with some minor modifications to the code, FP32 will be sufficient for the majority of the quantities of interest, up to and including turbulent stress budgets.
5 Transition to turbulence
We consider so-called K-type transition where a laminar baseflow as described by Schlatter et al.  [48, 49] is perturbed by one 2D and two oblique 3D Tollmien–Schlichting (TS) waves with amplitudes of $3\%$ and $0.1\%$ respectively (based on the laminar centerline velocity), all of which are individually stable. K-type (Klebanoff) transition refers to the instability being of fundamental type, i.e. the streamwise wavenumber of the primary and secondary instability are the same. The eigenvectors for the initial condition were computed using a Jupyter notebook [50] and superposed on top of the parabolic laminar Poiseuille flow. The Reynolds number is $Re_{b}=3333$ based on the constant bulk velocity, which corresponds to $Re_{cl}=5000$ based on the centerline velocity $U_{cl}$ of the initial parabolic velocity profile and the channel half-height $h$. The domain size is $5.61\delta\times 2.99\delta\times 2\delta$, adjusted to fit the chosen TS waves with $\alpha_{0}=1.12$ and $\beta_{0}=2.1$ as the streamwise and spanwise fundamental wavenumbers. For the compressible codes, instead of fixed mass flux, a constant pressure gradient forcing is applied in the streamwise direction to drive the flow. This would lead to a lower turbulent Reynolds number, but the initial growth of perturbations are only marginally affected. Thus, in all cases, a matching bulk Reynolds number of $Re_{b}=3333$ is maintained before turbulent breakdown.
We first focus on the expected behavior during transition, as illustrated in Figure 6. Panel a) shows the evolution of the two-dimensional (spanwise) modes $|\hat{u}(\alpha,\beta=0)|$, for integer $\alpha=0,1,\ldots$. It is always the maximum absolute value of the mode over the channel shown. It is clear that the mean-flow modes $\alpha=0$ and $\beta=0$ is only changing at $t>150h/U_{cl}$ corresponding to the establishment of the turbulent profile with a lower centerline component. The only other non-zero mode at $t=0$ is the 2D TS wave with $3\%$ energy. However, due to the nonlinearity of the flow and the triadic interactions, the flow quickly establishes a saturated 2D TS wave, with higher and higher 2D modes being energized, with a weak temporal decay.
Secondary instability, initiated by the $\beta=1$ modes, leads to a quick increase in the energy in all modes ($t>120h/U_{cl}$), the formation of characterisitic hairpin vortices ($t=136h/U_{cl}$, see Figure 7 )and subsequent breakdown to turbulence ($t>175h/U_{cl}$). The double-precision arithmetic allows us to resolve numerically all modes down to machine precision ($10^{-15}$) for Simson, but saturates at around $10^{-9}$ for the other solvers. Reducing to single precision increases the ambient noise level to about $10^{-8}$ for Simson, and around one order of magnitude higher for the other codes. Interestingly, there seems to be no interaction between these modes that would lead to a premature growth in the physically relevant modes. In contrast, similar studies using low-resolution simulations found a clear change of energy distribution and subsequent growth, which can be contained only using appropriate subgrid-scale models [48, 49]. It is noteworthy to highlight that not only the Fourier amplitudes and integral quantities are seemingly not affected by the lower precision, but also the actual vortical flow structures, as shown in Fig. 7. As opposed to lower-resolution simulations (as cited previously), turbulence is not appearing prematurely or disturbing the flow. The sharp gradients around the hairpin heads are well resolved without artifacts.
From Figure 8 we can conclude that the evolution of the individual modes, but also integral quantities such as the global friction or centerline velocities, is not dependent on the precision. For the rounding of the state and convective terms in the different solvers, we also observe that FP32 performs remarkably well, but when representing the state at lower-precision, the simulation becomes prone to stagnation (horizontal lines) or an immediate transition (State FP16). However, although the transitional case is sensitive, the amplitude of the initial conditions is still on the order of 0.10.10.10.11111%, and there will be a precision-dependent limit on the smallest disturbance amplitude the simulations would be able to capture. In addition, the geometry is a Cartesian channel, which motivates the study of a deformed geometry, such as the separating flow around a cylinder.
(a) Refer to caption (b) Refer to caption
Figure 6: Evolution of Fourier modes during K-type transition, obtained with Simson. The solid lines are FP64 (double precision), and the black dotted lines FP32 (single precision). a) shows the two-dimensional Fourier modes $|\hat{u}(\alpha,\beta=0)|$, starting with $\alpha=0$ on the top. (b) Three-dimensional Fourier modes $|\hat{u}(\alpha=1,\beta)|$ starting with $\beta=0$ on the top. The wall-normal maximum is shown in both cases.
Figure 7: Three-dimensional visualization of the flow right before breakdown to turbulence ($t=136$). Iso-contours of negative $\lambda_{2}=-0.03$ (scaled with channel half-width and center-line velocity) colored with the streamwise velocity. Only the lower channel-half is shown. Simulation performed with Simson using FP32 precision.
Figure 8: Evolution in time of amplitude of 2D modes $|\hat{u}(\alpha,\beta=0)|$ for transitional case. All simulations carried out in Neko and SSDC, except the reference case in Simson, shown in gray. Results from full FP32 and FP64 (a), State rounding (b), and rounding of the convective term (c) are shown.
Figure 9: Profiles for the cylinder at $Re_{D}=3900$ in Neko and SSDC with rounding the convective term, the state, and running the entire solver in single and double precision. The $C_{p}$ profile in the center of wake (a) and wake profile at four different locations in the wake $(0.58,1.06,1.54,2.02)$ (b), the blue-shaded interval is between time averages of two low-frequency modes as described by [51].
Figure 10: The Wasserstein distance between the probability distributions of the drag coefficient for the cylinder case for Neko and SSDC. A small distance means that the two distributions are similar.
Setup Avg. time ($D/U_{\infty}$) $f_{vs}$ $\phi_{s}$ $L_{r}$ $\bar{C}_{d}$ $-\bar{C}_{Pb}$
[52] 35 0.21 88 1.35 1.04 0.94
Neko
Full FP64 300.3 0.2097 86.62 1.48 0.9926 0.9159
Full FP32 319.1 0.2068 86.46 1.553 0.9911 0.8979
Convective FP32 100.0 0.2087 86.97 1.371 1.025 0.9496
Convective FP16 100.0 0.2087 86.62 1.481 1.004 0.9149
Convective bFloat16 100.0 0.2087 87.18 1.271 1.042 0.9775
Convective E4M3 73.13 0.204 87.49 1.182 1.057 1.003
Convective E5M2 77.46 0.2054 86.64 1.472 1.003 0.9148
State FP32 100.0 0.2087 86.8 1.402 1.016 0.9347
SSDC
Full FP64 300.0 0.2075 87.1 1.336 1.076 0.9731
Combined FP32 100.0 0.2035 86.7 1.391 1.058 0.9185
Combined FP16 100.0 0.2050 86.6 1.386 1.062 0.9475
Convective FP32 100.0 0.2050 87.5 1.149 1.117 1.0285
Convective FP16 100.0 0.2050 86.8 1.272 1.061 0.9233
Convective bfloat16 100.0 0.2050 86.6 1.405 1.054 0.9202
Convective E4M3 30.0 0.2099 87.1 1.205 1.069 0.9695
Convective E5M2 100.0 0.2064 86.8 1.272 1.093 0.9325
State FP32 100.0 0.2035 87.2 1.183 1.085 0.9756
State FP16 100.0 0.2099 86.8 1.386 1.063 0.9305
PadeLibs
Full FP64 69.0 0.2093 87.35 1.348 0.9932 0.9596
Convective FP16 36.6 0.2093 87.35 1.297 0.9932 0.9596
Convective E5M2 56.3 0.2097 88.86 1.028 1.073 1.0352
Table 4: Scalar values associated with the cylinder at $Re_{D}=3900$. Columns correspond to each setup name, the time statistics were collected for, the separation angle $\phi_{s}$, the recirculation length $L_{r}$, the drag coefficient $\bar{C}_{d}$, and the base pressure coefficient $\bar{C}_{Pb}$.
6 Separated flow — Cylinder at $Re_{D}=3900$
This section considers the flow around an infinite circular cylinder at $Re_{D}={U_{\infty}D}/{\nu}=3900$, where $D$ is the cylinder diameter and $U_{\infty}$ the free-stream velocity. We perform LES in PadeLibs, SSDC and Neko with approximately 512 grid points along the cylinder boundary, and a spanwise length of $2\pi D$ with $128$ grid points. There is extensive literature on this case, showing a significant spread in the simulation results [51]. The results are illustrated in Figure 9, which shows the velocity profiles in the wake and pressure distribution on the cylinder surface. Table 4 compares these results with the original LES by [52] and highlights wall quantities such as the drag coefficient and separation angle, as well as the length of the recirculation zone. Overall, the differences among the setups and solvers are comparable to the spread in the reference data. As such, for the simulations that do not diverge, this case indicates that other sources of uncertainty are more significant than the numerical precision when it comes to LES of separating flows such as the size of the domain and, in particular, averaging times.
Isolating the impact of lower precision might become clearer with longer averaging times, but due to the discrepancies among multiple reference data it is not clear whether the impact of precision can be isolated. A case with a stronger consensus among the reference data and where shorter averaging times are necessary would likely be better suited to evaluate only the impact of numerical precision.
For a more nuanced comparison, we employ the Wasserstein distance of the probability density functions of the drag coefficient $C_{D}$. The Wasserstein distance is a metric to compare the similarity between PDFs and has previously been employed for a similar purpose (impact of rounding and floating-point precision) in [6] for climate models where the similarity of the statistical description of the system is under consideration.
By comparing the Wasserstein distance between the drag coefficient $C_{D}$ for the different cases, the impact of averaging times is further amplified. We show the distance between the probability distributions of $C_{D}$ among the different runs in Figure 10. In this plot, the variability for the runs which were only carried out for 100 time units are clearly evident. For Neko, it is only for the longer averaging times that FP32 and FP64 consistently have a smaller Wasserstein distance than the poorer averaged results. While the convective FP32 might appear to differ, by comparing the longer FP64 run with exactly the same setup but averaged for a shorter time, we also see a large difference that is aligned with the low-frequency oscillations that are also described in [51].
7 Compressible flow around a wing section
[…]
This flow regime is characterized by a laminar separation bubble with transition and turbulent reattachment, as illustrated by plotting iso-contours of the second invariant of the velocity gradient tensor in Figure 11. The flow is initialized with a preliminary two-dimensional solution, the simulation is run in FP64 for $15$ convective time units ($C/U_{\infty}$), after which the statistics are computed for an extra $12$ convective time units for both considered precisions.
Table 5 compares the averaged integral loads and the separation bubble extension between the two SSDC runs and the reference data. Figure 12 shows the mean pressure coefficient and skin friction coefficient on the airfoil surface. Overall, single precision computations achieve similar results compared to double precision, and both compare well to the reference. Minor differences with respect to the reference are visible in the skin friction coefficient plot, close to the leading edge on the pressure side. These can likely be attributed to a slightly different discretization of the laminar boundary layer, which is very thin in this region. Figure 13 shows mean velocity profiles and the separation bubble along the suction side for both precisions. Both computations achieve very similar results even for these quantities. The time dependence of the separation is assessed considering the probability density function of the skin friction coefficient along the suction side of the airfoil, Figure 14. In particular, the same procedure as in [53], Fig. 13c, has been applied for e.g. binning. Even for this sensitive quantity, both solutions demonstrate remarkably similar results.
Figure 11: Laminar separation bubble with transition and turbulent reattachment observable from the iso-surfaces of the second invariant of the velocity gradient tensor ($Q=50$) colored by streamwise velocity ($u$).
Setup $t_{avg}U_{\infty}/C$ $C_{L}$ $C_{D}$ $C_{D_{P}}$ $C_{D_{sf}}$ $\text{x}_{1}|_{C_{f}=0}$ $\text{x}_{2}|_{C_{f}=0}$
Ref. [53] 7.7 0.621 0.0358 0.0220 0.0087 0.0999 0.6066
FP64 12 0.610 0.0355 0.0264 0.0090 0.1046 0.5965
Comb. FP32 12 0.614 0.0364 0.0275 0.0089 0.1004 0.6040
Table 5: Scalar aerodynamic results associated with the airfoil at $Re_{C}=50000$ and $M=0.5$. Columns correspond to each setup name, associated time-averaging duration (in $C/U_{\infty}units$), the lift coefficient $C_{L}$, the drag coefficient $C_{D}$ (split into pressure and skin friction components), and the start and end points of the separation bubble, $\text{x}_{1,2}|_{C_{f}=0}$.
(a) Mean pressure coefficient $\mathbf{C_{p}}$.
(b) Mean skin friction coefficient $\mathbf{C_{f}}$.
Figure 12: Mean $C_{p}$ and $C_{f}$ plotted along the surface of the airfoil. Black dashed line: reference data from [53]. Light blue line: double precision computation, Orange line: single precision computation with PCS.
Figure 13: Velocity profiles ($u(y)|_{x=x_{k}}$, continuous lines) along the airfoil at locations $x_{k}\in[0.1,1.0]$ and separation bubble ($u=0$, dot-dashed lines). FP64 in light blue, FP32 in orange.
Figure 14: Contour plot of the PDF of friction coefficient $C_{f}$ along airfoil chord. The PDF is scaled given a non-uniform bin size along $x$, more details can be found in [53] (Eq. 3.1, Fig. 13).
8 Practical experiences
This section summarizes our experiences in running our solvers natively in FP32. We comment on both performance gains (particularly on GPUs) and a range of issues that needed to be addressed for the solvers to run problem-free. While the identified problems differ among the solvers, collectively they serve as useful pointers to what should be treated with extra care when implementing an FP32-capable solver.
8.1 Performance improvements with FP32 over FP64
One of the most important practical outcomes from the simulation campaign was evaluating the performance impact of using FP32 instead of FP64 for the different test cases. Recall that F32 was the only precision for which the rounding was available directly on the hardware and not emulated, thus for the other precisions such measurements were not possible to obtain.
We gained most experience from our runs with Neko where a performance boost around 2×2$\times$2 × was expected as most kernels operate in the memory-bound domain. However, it was observed that the performance improvement changed depending on whether server-grade or consumer GPUs were used, as illustrated in Table 6.
In general, it was found that a larger problem size was needed to obtain 2×2$\times$2 × performance improvement on the server grade GPUs going from FP64 to FP32. For the cylinder case, for example, running on the AMD Instinct MI250X GPUs the performance was more or less exactly 2×2$\times$2 × faster with single precision. However, when a smaller turbulent channel flow case was considered this was no longer as evident as the problem size was insufficient to hide the latency of kernel launches and oversubscribe the available computational resources.
A comparison between the server-grade Nvidia A100 and the consumer-level Nvidia RTX 4080 revealed that, for single precision runs, the A100 achieved a performance increase of approximately 1.5×1.5$\times$1.5 × over its double precision counterpart in the smaller channel flow case. In contrast, the RTX 4080 demonstrated more than double the performance in single precision, and its FP32 performance matched that of the much more expensive A100 operating in FP64. This suggests that it is feasible to perform DNS on consumer-grade GPUs using FP32, achieving comparable performance to traditional FP64 simulations on high-end server GPUs. In this case, simulations executed in FP32 on the RTX 4080 produced results equivalent to those previously obtained with FP64 on the A100—within the same runtime.
GPU $\beta$ $\pi_{\text{FP}64}$ $\pi_{\text{FP}32}$ FP32 vs FP64
RTX4080 0.72 TB/s 0.76 TFlop/s 48.7 TFlop/s 22.5×2-2.5$\times$2 - 2.5 ×
A40 0.7 0.58 37.4 22.5×2-2.5$\times$2 - 2.5 ×
A100 1.56 9.7 19.4 1.52×1.5-2$\times$1.5 - 2 ×
Table 6: Bandwidth to global memory (DRAM or HBM) $\beta$ in TB/s, and performance in TFlop/s
8.2 Stagnation
As previously mentioned, the issue of stagnation can become prevalent when using lower precision when summing a large and small number.
The most straightforward and easy to address impact of lower arithmetic precision was observed for example in longer simulations (integration times of $\mathcal{O}(10^{3})$ convective units or larger) with smaller time steps. This was observed with precisions as high as FP32 and as issues with correct estimation of the simulation time (usually calculated as $t_{\rm new}=t_{\rm old}+\Delta t$) as well as some time-dependent if statements (such as those that control the calculation of additional variables at constant intervals). While this was observed mainly in Simson on the longest simulations (integration times of $10000$ convective times or higher) when compiled with FP32, it can easily happen for other codes, especially for smaller time steps. This suggests that integration times $t$ and its related quantities (weights etc.) should use higher precisions (FP64), even when the code is compiled with lower precision.
Another similar observation was the impact of precision on runtime collection of solution statistics. This is related to the addition of the new sample to the previous set, usually done as a variant of $S_{\rm new}=wS_{\rm old}+(1-w)s$, where $S_{\rm new}$ and $S_{\rm old}$ are new and old statistics and $s$ is the new sample. The weighting parameter here, $w=t/(t+\delta t)$, is again highly impacted by precision for short sampling times $\delta t$ (i.e., frequent sampling) and long integration times $t$. This was, for instance, observed as non-zero residuals in the transport equation of Reynolds stresses when calculated from runtime statistics, while not observed when calculated from a few hundred snapshots, despite having far fewer samples.
It was also relevant when computing larger dot-products in Neko, where a naive implementation can run into stagnation issues. This can be remedied through tree-reductions or performing the accumulation in higher precision. This again suggests that to ensure robustness and accuracy of runtime statistics, such calculations and accumulation should be performed in FP64.
8.3 Arithmetic errors
Among all the computations performed with SSDC 3.1.3, the TS-wave case has shown some implementation details that can be useful in practice. The default FP64 computation is used as a reference, considering the emulated precision with the CPFloat library on both the state and operator terms, reasonable results are obtained as previously discussed. However, when compiling the code in single precision (which is not what the implementation is originally designed for), the computation exhibits nonphysical artifacts in the flow, leading to larger numerical noise and premature transition to turbulence (see Figure 15). Upon close examination, we could trace the origin of the error to the convective term, specifically in the form of a loss of freestream preservation. In simpler terms, the computation of geometric quantities (Jacobians, cell volumes and normals) in single precision resulted in a non-watertight grid. The correct behavior was recovered when switching these computations back to FP64. Similar observations could also be seen when computing the Lambda-2 criterion in Neko, where a naive way of obtaining the required eigenvalues was sensitive to the lower floating point precision. The most straightforward way to remedy this was by making parts of this computation in FP64, as it was neither time nor memory critical.
Figure 15: Evolution in time of the amplitude of 2D modes $|\hat{u}(\alpha,\beta=0)|$ for the transitional case. All simulations are carried out in SSDC. Results from full FP64 and FP32, FP32 emulated with PCS, and compiled FP32 except for the computation of metric terms (MT) in FP64 are shown.
For Simson, it was also evident that in the Chebyshev discretization, points are highly clustered near the ends of the interval, and the derivative matrices are fully populated. As a result, the derivative at any given point is calculated as a weighted sum of values from all other points, with both the weights and function values varying significantly across the domain. A quick examination of the differentiation matrix for first and second derivatives using 257 points (the grid used for the channel flow at $Re_{\tau}\approx 550$) reveals differences of approximately five orders of magnitude between terms in the same row (terms multiplied by values on wall-normal grid points), with the largest discrepancies occurring near the walls. This effect is further amplified in a velocity-vorticity formulation, where higher-order derivatives (up to the fourth order) are required. These factors suggest a high sensitivity to numerical precision, particularly near the wall, which is likely the root cause of the observed issues with the pressure solver in Simson (Section 4.1). Note that the employed Chebyshev-tau method is less sensitive to the poor conditioning.
Similarly, computing the streamwise and spanwise velocity components requires reconstructing values from wall-normal velocity and vorticity (with varying magnitudes) and solving a Poisson equation for the mean components. Both processes can be highly sensitive to numerical precision, which explains the sensitivity of higher moments of streamwise velocity discussed in Section 4.1.
It is important to highlight that while precision was expected to significantly influence derivatives and flow dynamics, interestingly and somewhat unexpectedly, these sensitivities did not result in irrecoverable errors. In other words, many of the issues were found to be resolved by performing as little as a single time step in FP64.
9 Conclusions
The main outcome of this work is the demonstrated strong evidence that high-fidelity simulations of wall-bounded turbulence, including direct numerical and large-eddy simulation (DNS and LES), do not necessarily require FP64 arithmetic and can be performed at lower precisions, such as FP32, with minimal impact on the results. We show that this is possible not only for simpler canonical flow cases such as channels, but also for more complex flows exhibiting separation and transition. Our results are consistent across different formulations of governing equations considered in the paper, as well as underlying discretization methods and their implementations in different codes. Essentially perfect overlap is obtained between FP64 and FP32 profiles of various quantities of interest, including, for example, turbulent kinetic energy budgets and high-order statistics. Although we do not focus on performance analysis, we show that using FP32 significant acceleration, up to the ideal 2×2$\times$2 × for a well-optimized code is possible compared to standard FP64. In particular, in FP32 a DNS of selected cases can be run on a single consumer GPU instead of more expensive server-grade cards. In addition, the performance improvement is not limited to computational time but also translates to reduced storage demands. Considering that large-scale CFD simulations utilize hundreds of millions of core hours yearly, all codes should apply significant effort to utilize lower-precision arithmetic, thus saving energy, storage, and money for the same scientific outcomes. However, some caution must be exercised when porting existing solvers, as one likely needs to retain some of the operations in FP64 (see Section 8).
In addition, the paper explores using even lower precisions by means of software emulation. The results are encouraging, with sometimes very low precision, like E5M2, providing good results when applied to only the convective term. A natural way forward is to investigate the native implementation of lower precision formats that are supported by modern hardware. Success is likely conditioned on using mixed precision with the concrete realization of the latter tightly coupled with the numerical method used by the code.
On a subjective note, the results of this study stand in quite strong contradiction with our initial expectations. We anticipated to quickly find a (high) precision threshold, after which the results become unusable. In particular, for flows with transition, we were very skeptical about the possibility of using anything but FP64. However, we were proved wrong and our hope is that this paper will motivate an equally skeptical reader to give reduced precision a chance.
More generally, the question of using arithmetic precisions lower than FP64 should not be viewed as one with a binary yes or no answer, but instead as another hyper-parameter in a simulation. In large-scale simulations, it is often the norm to assess the influence of multiple parameters before the production run starts. These parameters usually include things such as grid resolution, time-step size (i.e. Courant number), domain size, residual thresholds of iterative solvers, and so on. In our view, precision should also become one such parameter, even though the control over it is usually quite rough (i.e., a jump from FP64 to FP32 instead of a smooth transition). In fact, one could argue that even in cases where reduced precision leads to (small) discrepancies in the results, it should still be viewed as just another source of uncertainty in the result, albeit a bias. In that sense, viewing a computational simulation as an optimization problem to maximize the accuracy of the output for a given cost, arithmetic precision must be balanced against other sources of uncertainty such as those related to time averaging, residual, or resolution. Similarly, in parametric studies and simulation campaigns a lower cost per simulation would enable additional simulations, which again should be balanced again other parameters for an objective function of maximizing the knowledge gained from the campaign for a given computational cost.
Despite the wider scope of this study compared to the available literature, it is important to keep in mind that it is still fairly limited compared to the wider applications of high-fidelity simulations. For example, Reynolds numbers of this study were still relatively low and complex geometries or physics were avoided. Future extensions of this work could benefit from a wider variety of test cases, including supersonic and hypersonic flows, combustion, and complex geometries, among other aspects.
The other shortcoming of this work was the lack of a quantitative and theoretical measure for the effect of precision in different regions of the domain. While developing such measures is relatively easy for compressible flows with explicit time stepping and no iterative solvers, it was proved difficult for an incompressible solver such as Neko with a fractional step algorithm, coarse grid solvers, and iterative methods. Developing a metric to quantify the impact of precision on the governing equations and the solution will be another topic for future research.
Acknowledgments
This project was initiated during the CTR Summer Program 2025 in Stanford [32], for which we gratefully acknowledge the financial support to be able to participate. Computer time was provided by the National Infrastructure for Computing in Sweden (NAISS). The authors gratefully acknowledge the HPC resources provided for this collaborative effort by the Supercomputing Laboratory at King Abdullah University of Science & Technology (KAUST) in Thuwal, Saudi Arabia. The authors gratefully acknowledge the scientific support and HPC resources provided by the Erlangen National High Performance Computing Center (NHR@FAU) of the Friedrich-Alexander-Universität Erlangen-Nürnberg (FAU). The hardware is funded by the German Research Foundation (DFG). This project has received funding from the European High-Performance Computing Joint Undertaking (JU) under grant agreement No 101092621. The JU receives support from the European Union’s Horizon Europe research and innovation programme and Germany, Italy, Slovenia, Spain, Sweden and France.
Appendix A Rounding configurations
Table 7 lists the rounding performed with each code for all the test cases.
CFD Code Format Rounding type Fully developed turbulence Transition to turbulence Separated flow Cylinder Compressible flow Airfoil
Neko FP64 Full
FP32 Full
Convective
State
FP16 Full
Convective
State
bfloat16 Convective
State
E4M3 Convective
State
E5M2 Convective
State
PadeLibs FP64 Full
FP32 Full
Convective
State
FP16 Full
Convective
State
bfloat16 Convective
State
E4M3 Convective
State
E5M2 Convective
State
Simson FP64 Full
FP32 Full
Convective
State
FP16 Full
Convective
State
bfloat16 Convective
State
E4M3 Convective
State
E5M2 Convective
State
SSDC FP64 Full
FP32 Full
Combined
State
Operator
FP16 Full
Combined
State
Operator
bfloat16 Combined
State
Operator
E4M3 Combined
State
Operator
E5M2 Combined
State
Operator
Table 7: Solvers with different roundings for each test case.